OpenStudy (anonymous):

METAL WILL BE AWARDED Write down the next three rows to continue this sequence of equations. 1 = 1 = 1^3 3 + 5 = 8 = 2^3 7 + 9 + 11 = 27 = 3^3 13 + 15 + 17 + 19 = 64 = 4^3

hero (hero):

Well we at least know that based on the given pattern, it is obvious that the next three rows will be 5^3 6^3 7^3 right?

OpenStudy (anonymous):

Yes but there is more

hero (hero):

Of course. The next step is so evaluate 5^3, 6^3, and 7^3

hero (hero):

In other words, find the integer values of each.

OpenStudy (anonymous):

I think there's more to solve it like the addition to it

hero (hero):

Yes, but @LeadSkills, do you know how Open Study works? The goal here is to "guide" not just give the answer all at once.

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

I am thinking of this problem and I see a pattern with the numbers and the exponents

hero (hero):

Great. Would you mind posting what you believe the next three rows are?

OpenStudy (anonymous):

Well numbers are going down obvious so the 1st one is adding 4

hero (hero):

Hint: Notice that the sums on each row are consecutive odd integers: {1, 3, 5, 7, 9, 11, 13, 15, 17, 19...}

OpenStudy (anonymous):

Now I understand

hero (hero):

The best way to demonstrate that you understand is to post what you believe are the next three rows.

OpenStudy (anonymous):

Okay I will Hero

OpenStudy (anonymous):

1 = 1 = 13 3 + 5 = 8 = 23 7 + 9 + 11 = 27 = 33 13 + 15 + 17 + 19 = 64 = 43 21 + 23 + 25 + 27 + 29 = 125 = 53 31 + 33 + 35+ 37 + 39 + 41 = 216 = 63 43+ 45 + 47 + 49+ 51+53 + 55 = 343 = 73

OpenStudy (anonymous):

@Hero 1 = 1 = 13 3 + 5 = 8 = 23 7 + 9 + 11 = 27 = 33 13 + 15 + 17 + 19 = 64 = 43 21 + 23 + 25 + 27 + 29 = 125 = 53 31 + 33 + 35+ 37 + 39 + 41 = 216 = 63 43+ 45 + 47 + 49+ 51+53 + 55 = 343 = 73

hero (hero):

I believe you have it. The only discrepancy seems to be that you left out the carets for 1^3, 2^3, 3^3, etc...

OpenStudy (anonymous):

Ye It didnt transfer in open study

hero (hero):

Great job though.

OpenStudy (anonymous):

I have and other question to solve and I want to try to solve it.

OpenStudy (anonymous):

A train can hold 78 passengers. The train starts out empty and picks up 1 passenger at the first stop, 2 passengers and the second stop, 3 passengers at the third stop, and so forth. After how many stops will the train be full?

OpenStudy (anonymous):

Hold on